OpenStudy (anonymous):

Solve the equation one-half times x plus three-fourths equals one-fourth for x.

OpenStudy (anonymous):

can someone plz help

OpenStudy (anonymous):

@stonewoods can you please help me

OpenStudy (anonymous):

hold on one moment love ^~^

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

x=-1, i worked it out on paper; i hope that helps!

OpenStudy (anonymous):

yes can you help me with some more please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a.

OpenStudy (anonymous):

i think its 1/5

OpenStudy (anonymous):

well this are the answer that come with it a equals negative seven-fortieths a equals negative seven-tenths a equals negative twenty-one fortieths a equals negative twenty-one tenths

OpenStudy (anonymous):

im not sure, sorry :/

jigglypuff314 (jigglypuff314):

\[\frac{ 1 }{ 2 }x + \frac{ 3 }{ 4 } = \frac{ 1 }{ 4 }\]multiply through by 4 to get 2x + 3 = 1

OpenStudy (anonymous):

thank you @jigglypuff314 can you help me with this one Solve negative two-thirds times a plus two-fifths equals three-fourths for a.

jigglypuff314 (jigglypuff314):

0.o erm, get common denominator... I'ma try 60 so... ... -40a + 24 = 45